Use the following information to answer the question: Qs = 4/5 P Qd = 20 - 4/5 P What is the amount of tax revenue if a $2 tax is imposed on each item sold? $32.40 $3 $18.4 $36 $13.50
Answer #1

Qs = 4/5 P

Qd = 20 - 4/5 P

Since a tax of $2 is imposed, the burden of the tax falls on the sellers. The price at which the sellers will sell will now be higher.

Thus, the supply equation now becomes Qs = 4/5 (P + 2)

Equating the demand and supply equations,

Qs = Qd

4/5 (P + 2) = 20 - 4/5 P

8/5 (P + 1) = 20

P = $11.5

At this price the quantity supplied is Qs = 4/5 P = 4/5 (11.5) = 9.2

=> Qs = 9.2

Note: We have calculated the quantity supplied using the old supply equation because the effect of the tax has already been taken into consideration while calculating equilibrium price. Thus, the price of $11.5 is the 'P' after tax that we use here.

Tax revenue = Tax x Quantity (that the firms stand ready to supply)

Tax revenue = 2 x 9.2

Tax revenue = $18.4

Correct answer : c) $18.4